# Gravitational wave propagation in hybrid metric-Palatini gravity

### v1

@@ VI Conclusions
In this work we have analyzed the propagation modes and polarization states of gravitational waves in the hybrid metric-Palatini theory of gravity in its scalar-tensor representation both from the linearization and Newman-Penrose formalisms. Furthermore, we have analyzed the results in light of recent outcomes from a cosmological perspective Rosa:2024pzo , in an attempt to qualify the falsifiability of the theory.
@@
In the linearization formalism, we found that the theory exhibits massless tensor modes that uncouple from the other degrees of freedom in the transverse-traceless gauge, and two generally massive scalar modes that decouple from each other, but their intertwined relation corresponds to a single degree of freedom, despite being able to excite up to two scalar polarization states, namely the longitudinal and the breathing modes. We found that the masses of the scalar modes depend on the background values of the potential $V$ and its partial derivatives, implying that one can always fine-tune the potential to eliminate the masses of the scalar modes and obtain a speed of propagation equal to the speed of light.
@@
Moreover, to account for the massive scalar modes, we have resorted to an adaptation of the Eardley-Lee-Lightman formalism which is built from the Newman-Penrose formalism, since the former originally was valid only for nearly null plane gravitational waves. Through this analysis, we have found that, at the linear level, the hybrid metric-Palatini gravity presents the same tensor modes as GR, although the $+$ polarization mode exhibits a sort of squeezing behavior, and that the breathing and longitudinal modes are excited by a single degree of freedom, in agreement with the results from the linearization method. We recall that in modified gravity theories the 00-component of the perturbed metric field needs not to be nonvanishing as it is encompassed in the Bardeen gauge invariant scalars and is associated with the longitudinal polarisation mode Dong:2025ddi , despite the TT-gauge being applied to the $ij$-components subspace.
@@
In general, in a cosmological spacetime given by the Friedmann-Robertson-Walker, the dispersion relation for the tensor modes can be parametrised by $h_{A}^{\prime\prime}+\left(2+\bar{\nu}+\nu_{A}\right)\mathcal{H}h_{A}^{\prime}+\left(1+\bar{\mu}+\mu_{A}\right)k^{2}h_{A}=0$, where the index $A\in\{L,R\}$, standing for the left-hand or right-hand circular polarisation modes, respectively, (which are built from the plus and cross tensor modes as $h_{A}=\frac{1}{\sqrt{2}}\left(h_{+}\pm ih_{\times}\right)$, the primes denote derivatives relatively to the conformal time, $\mathcal{H}$ is the Hubble parameter in conformal time, and the parameters $\nu_{A},~{}\mu_{A}$ label the effects of the parity violation which include the amplitude and velocity birefringences of the GWs, whilst the parameters $\bar{\nu},~{}\bar{\mu}$ arise from other modifications such as frequency-independent effects that modify the GWs speed and friction, as well as Lorentz-violating terms as frequency-dependent damping and nonlinear dispersion of GWs ParityViolation1 ; ParityViolation2 . Moreover, the parameters $\bar{\mu},~{}\mu_{A}$ are associated with the speed of the gravitational wave, whilst $\bar{\nu},~{}\nu_{A}$ provides an amplitude modulation to the gravitational waveform. Therefore, we find no amplitude birefringence, as we have not considered a cosmological spacetime on which the GW propagates (we leave this discussion for future work), and concerning possible terms associated with velocity birefringence, which currently are unconstrained by the data Macarena , we do not found any as the dispersion relation of the auxiliary field, $\chi_{\mu\nu}$, is the same as the original perturbed metric field, $\delta g_{\mu\nu}$, and is identical to the GR result. Therefore, the origin of the ”squeezing” feature that appears in the non-linear regime cannot be identified by the birefringence approach from the linearisation method, at least considering a Minkowskian background far from the source.
@@
To analyze the observational implications of the theory, we have constrained the scalar fields and the scalar potential following recent results obtained through a dynamical system analysis of cosmology and a compatibility with solar system dynamics. Furthermore, we assumed the waves to be approximately planar due to the large distance to their sources. Under these constraints, the scalar fields are exponentially suppressed at late times, consequently causing a suppression of the scalar modes at present times and reducing the differences between the tensor modes to the previously mentioned squeezing behavior in the $+$ polarization.
@@
These results imply that, for the theory to be simultaneously acceptable from both cosmology and gravitational wave physics perspectives under these assumptions, the squeezing effect is the only possible differentiation with respect to GR. However, because GR is included as a particular limit of the hybrid metric-Palatini gravity, the absence of a detection of the squeezing effect renders the hybrid metric-Palatini gravity unfalsifiable from the point of view of the propagation of gravitational waves. The same conclusion can be traced about the distinction between GR and the metric and Palatini $f(R)$ models, by taking the limits $\varphi=0$ and $\psi=0$, respectively.
@@
The hybrid metric-Palatini theory encompasses other gravity models, such as the metric and Palatini $f(R)$ models, as well as non-linear combinations of both. Our results are compatible with the polarization and degrees of freedom found in the literature for those. Moreover, to assess the different polarization states or behaviors, future gravitational wave detectors that may allow for different angles between pairs of interferometers would be required. In fact, current differential-arm detectors, also known as quadrupolar antennas, such as LIGO and Virgo, are less sensitive to scalar modes than to tensor ones (despite this varies with the sky location of the source although not very sharply), and fail to distinguish between the longitudinal and the breathing polarization modes LIGO:technicalnote . We also note that in massive scalar-tensor theories, the longitudinal polarization mode is suppressed relative to the breathing one by a factor $(\lambda/\lambda_{C})^{2}$, being $\lambda$ the wavelength of the gravitational wave, and $\lambda_{C}$ the Compton wavelength of the massive scalar CMWill . Nevertheless, attempts to enhance the scalar detection have been proposed Maggiore:2020 .
@@ Acknowledgements.
We thank Nicola Tamanini, Maxence Corman, Macarena Lagos, David Hilditch, Franscisco Lobo, Lara Sousa, and José Fonseca for fruitful discussions. C.G. acknowledges Fundação para a Ciência e Tecnologia (FCT) through Project UID/04650 - Centro de Física das Universidades do Minho e do Porto, and Project 2024.00252.CERN, and European Commission through COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology). MASP acknowledges support from the FCT through the Fellowship UI/BD/154479/2022, through the Research Grants UIDB/04434/2020 and UIDP/04434/2020, and through the project with reference PTDC/FIS-AST/0054/2021 (“BEYond LAmbda”). MASP also acknowledges support from the MICINU through the project with reference PID2023-149560NB-C21.

### v2

@@ VI Conclusions
In this work we have analyzed the propagation modes and polarization states of gravitational waves in the GHMPG in its scalar-tensor representation both from the linearization and Newman-Penrose formalisms. Furthermore, we have analyzed the results in light of recent outcomes from a cosmological perspective Rosa:2024pzo , in an attempt to qualify the falsifiability of the theory.
@@
In the linearization formalism, we found that the theory exhibits massless tensor modes that uncouple from the other degrees of freedom in the TT-gauge, and two generally massive scalar modes that decouple from each other, but their intertwined relation corresponds to a single degree of freedom, despite being able to excite up to two scalar polarization states, namely the longitudinal and the breathing modes. We found that the masses of the scalar modes depend on the background values of the potential $V$ and its partial derivatives, implying that one can always fine-tune the potential to eliminate the masses of the scalar modes and obtain a speed of propagation equal to the speed of light.
@@
Moreover, to account for the massive scalar modes, we have resorted to an adaptation of the Eardley-Lee-Lightman formalism which is built from the Newman-Penrose formalism, since the former originally was valid only for nearly null plane gravitational waves. Through this analysis, we have found that, at the linear level, the GHMPG presents the same tensor modes as GR, although the ++ polarization mode exhibits a sort of squeezing behavior, and that the breathing and longitudinal modes are excited by a single degree of freedom, in agreement with the results from the linearization method. We recall that in modified gravity theories the 00-component of the perturbed metric field needs not to be nonvanishing as it is encompassed in the Bardeen gauge invariant scalars and is associated with the longitudinal polarization mode Dong:2025ddi , despite the TT-gauge being applied to the $ij$-components subspace.
@@
In general, in a cosmological spacetime given by the Friedmann-Robertson-Walker metric, the dispersion relation for the tensor modes can be parameterized by $h_{A}^{\prime\prime}+\left(2+\bar{\nu}+\nu_{A}\right)\mathcal{H}h_{A}^{\prime}+\left(1+\bar{\mu}+\mu_{A}\right)k^{2}h_{A}=0$, where the index $A\in\{L,R\}$, standing for the left-hand or right-hand circular polarization modes, respectively, (which are built from the plus and cross tensor modes as $h_{A}=\frac{1}{\sqrt{2}}\left(h_{+}\pm ih_{\times}\right)$, the primes denote derivatives relatively to the conformal time, $\mathcal{H}$ is the Hubble parameter in conformal time, and the parameters $\nu_{A},~\mu_{A}$ label the effects of the parity violation which include the amplitude and velocity birefringences of the GWs, whilst the parameters $\bar{\nu},~\bar{\mu}$ arise from other modifications such as frequency-independent effects that modify the GWs speed and friction, as well as Lorentz-violating terms as frequency-dependent damping and nonlinear dispersion of GWs ParityViolation1 ; ParityViolation2 . Moreover, the parameters $\bar{\mu},~\mu_{A}$ are associated with the speed of the gravitational wave, whilst $\bar{\nu},~\nu_{A}$ provides an amplitude modulation to the gravitational waveform. Therefore, we find no amplitude birefringence, as we have not considered a cosmological spacetime on which the GW propagates (we leave this discussion for future work), and concerning possible terms associated with velocity birefringence, which currently are unconstrained by the data Macarena , we do not found any as the dispersion relation of the auxiliary field, $\chi_{\mu\nu}$, is the same as the original perturbed metric field, $\delta g_{\mu\nu}$, and is identical to the GR result. Therefore, the origin of the ”squeezing” feature that appears in the non-linear regime cannot be identified by the birefringence approach from the linearization method, at least considering a Minkowskian background far from the source.
@@
To analyze the observational implications of the theory, we have constrained the scalar fields and the scalar potential following recent results obtained through a dynamical system analysis of cosmology and a compatibility with solar system dynamics. Furthermore, we assumed the waves to be approximately planar due to the large distance to their sources. Under these constraints, the scalar fields are exponentially suppressed at late times, consequently causing a suppression of the scalar modes at present times and reducing the differences between the tensor modes to the previously mentioned squeezing behavior in the ++ polarization.
@@
These results imply that, for the theory to be simultaneously acceptable from both cosmology and gravitational wave physics perspectives under these assumptions, the squeezing effect is the only possible differentiation with respect to GR. However, because GR is included as a particular limit of the GHMPG, the absence of a detection of the squeezing effect renders the GHMPG unfalsifiable from the point of view of the propagation of gravitational waves. The same conclusion can be traced about the distinction between GR and the metric and Palatini $f(R)$ models, by taking the limits $\varphi=0$ and $\psi=0$, respectively.
Moreover, we should stress that our analysis finds the highest number of polarization states that can exist in the GHMPG model. In fact, some subclasses of models that obey the observational constraints that we used can have fewer modes. For instance, the $f(R)=R^{2}$ model has no propagating scalar, vector, or tensor mode whatsoever Alvarez-Gaume:2015rwa . Nevertheless, there are more $f(R)$ models which are also degenerate in that sense of absence of degrees of freedom or instability, both at the linearized level on top of a maximally-symmetric background (thus, including Minkowski) Casado-Turrion:2024esi . Such systematization is not easily extended to a two-variable function, such as in the GHMPG. Moreover, there are complementary classification systems that resort to NP-quantities, which are the Petrov-Pirani-Penrose Petrov:2000bs and the Plebanski classifications McIntosh:1981 . These classification systems are beyond the scope of the present work.
@@
The GHMPG theory encompasses other gravity models, such as the metric and Palatini $f(R)$ models, as well as non-linear combinations of both. Our results are compatible with the polarization and degrees of freedom found in the literature for those. Moreover, to assess the different polarization states or behaviors, future gravitational wave detectors that may allow for different angles between pairs of interferometers would be required. In fact, current differential-arm detectors, also known as quadrupolar antennas, such as LIGO and Virgo, are less sensitive to scalar modes than to tensor ones (despite this varies with the sky location of the source although not very sharply), and fail to distinguish between the longitudinal and the breathing polarization modes LIGO:technicalnote . We also note that in massive scalar-tensor theories, the longitudinal polarization mode is suppressed relative to the breathing one by a factor $(\lambda/\lambda_{C})^{2}$, being $\lambda$ the wavelength of the gravitational wave, and $\lambda_{C}$ the Compton wavelength of the massive scalar CMWill . Nevertheless, attempts to enhance the scalar detection have been proposed Maggiore:2020 .
@@ Acknowledgements.
We thank Nicola Tamanini, Maxence Corman, Macarena Lagos, David Hilditch, Francisco Lobo, Lara Sousa, and José Fonseca for fruitful discussions. C.G. acknowledges Fundação para a Ciência e Tecnologia (FCT) through Project UID/04650 - Centro de Física das Universidades do Minho e do Porto, and Project 2024.00252.CERN. JLR is supported by the Project PID2022-138607NBI00, funded by MICIU/AEI/10.13039/501100011033 (“ERDF A way of making Europe” and “PGC Generación de Conocimiento”). MASP acknowledges support from the FCT through the Fellowship UI/BD/154479/2022, through the Research Grants UIDB/04434/2020 and UIDP/04434/2020, and through the project with reference PTDC/FIS-AST/0054/2021 (“BEYond LAmbda”). MASP also acknowledges support from the MICINU through the project with reference PID2023-149560NB-C21. C.G. and MASP acknowledge the European Commission through COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology).